OpenStudy (anonymous):

Write the equation of the line that passes through (-1, 4) with a slope of -1 in standard form.

OpenStudy (anonymous):

I'm studying this now so I'm not sure if this is right. If you plug the numbers in point slope form : y-y1=m(x-x2) then y-4=-19x-4) so add four to both sides and get y=-1x+8. If you reorder it you should get x+y=8. Can anyone check to make sure this is right?

OpenStudy (anonymous):

I actually have four options. -x + 4y = -1 x - 4y = -1 x + y = 3 x - y = -3

OpenStudy (anonymous):

Wow. I am so great at math. -_- I don't know I can try again

OpenStudy (anonymous):

I'd do it like this: y-4=-(x+1). Distribute the -, so: y-4=-x-1. Then, move the x to the other side, and add the 4 to the other side, and you have: x+y=3.

OpenStudy (anonymous):

Ooooh, alright! Thank you guys! (:

OpenStudy (anonymous):

Yup that makes sense^ I always have dumb mistakes xD

OpenStudy (anonymous):

It's ok. Don't be hard on yourself @__mia_rachel__